OpenStudy (anonymous):

Someone please help me!! I will medal!! Integrate 1/((x^2 +2)^2)

OpenStudy (anonymous):

How do I expand this using partial fractions?

OpenStudy (anonymous):

I don't need the whole answer

OpenStudy (freckles):

I don't think partial fractions would work here because that looks like this \[\frac{Ax+B}{x^2+2}+\frac{Cx+D}{(x^2+2)^2}\] all the constants should be 0 except D unless I'm mistaken

OpenStudy (freckles):

I think you might want to do a trig sub

OpenStudy (rational):

^

OpenStudy (anonymous):

Oh really? It shows up in my partial frac notes package....

OpenStudy (freckles):

I think we can do it over the complex numbers

OpenStudy (anonymous):

1/x4 + 4x² + 4

OpenStudy (anonymous):

²

OpenStudy (anonymous):

why won't ln work for it?

OpenStudy (anonymous):

i guess ill just use trig sub... thanks anyways guys!!!

OpenStudy (rational):

there is a nice partial fraction way of working this if you allow some guessing

OpenStudy (rational):

nvm it got complicated :/ we better stick to trig substitution \(x = \sqrt{2}\tan u\)